For the equation CH4 + 2O2 →CO2 + H2O + 803.1 kJ, how much heat will be released if 20 g of CH4 reacts with sufficient O2? _______

A. 883 kJ
B. 883 kJ/mol
C. 1,004 kJ
D. 1,004 kJ/mol
E. 1.004 kJ
1
Expert's answer
2018-06-29T05:02:35-0400
n(CH4) = m/M = 20g/16g/mol = 1.25 mol;

∆H' = 1.25*∆H = 1.25*803.1 = 1,004 kJ.

Answer: C. 1,004 kJ